# TOP-Nav: Legged Navigation Integrating Terrain, Obstacle and Proprioception Estimation

### v2

@@ IV-B Integrated Path Planner
IV-C Proprioception Advisor
Figure 3: We demonstrate the variation of the normalized ${c_{t}}$ when the robot traverses on different types of terrain. The traversal on bush reveals unstable postures and a failure to track velocity commands, resulting in a decrease in $c_{t}$. Conversely, navigating on pavement exhibits a stable $c_{t}$.
We design the proprioception advisor to identify motion abnormalities that may arise from unexpected external disturbances. A straightforward approach is to use the velocity tracking error [11], but its effectiveness can be hindered by the noisy velocity estimation on real hardware. Considering the interactions between the robot and its environment, the ideal metrics should be both observable and comprehensive. To address this, we utilize the value function $c_{t}$ estimated from $\pi_{\textrm{critic}}$, which synthesizes multiple metrics and provides more stable results by incorporating historical observations.
The original $\pi_{\textrm{critic}}$ requires privileged observations, we train another value function estimation network using observable motion states for the proprioception advisor to be deployed on hardware. The estimated motion evaluation $c_{t}$ is normalized using a sigmoid function. As demonstarted in Fig. 3, the motion evaluation will decline sharply when the robot encounters a locomotion failure, potentially due to transitioning onto challenging terrain. Implementation details for the reward design and normalization are provided in the appendix.
Within the path planner, we first directly utilize the motion evaluations to enhance the robot’s awareness of unforeseen disturbances. This involves estimating collisions ($\bm{M}_{P}$) beyond the obstacle estimator and integrating them into the combined cost map ($\bm{M}_{C}$). Previous works generate fix-sized virtual obstalces when proprioception feedback reaches a certain threshold [14], the design overlooks the variation of motion evaluations within each interval. We propose an continuously varied proprioception cost along the y-axis in $\bm{M}_{P}$.
here $k_{P}$ is hyperparameter. Since a lower motion evaluation indicates a potential challenging terrain or an invisible obstacle in the current direction of the robot, we allocate $c_{t}$ to the centriod column in ${\bm{M}_{P}}$ and decrease it towards the edges.
IV-D Vision-based Terrain Estimation
The proposed terrain estimator incorporates terrain traversability into the path planner. Firstly, we construct a mapping from visual observations to reference terrain traversabilities using prior knowledge.
Given that terrains like stairs, slopes, or discrete small steps can be effectively identified from depth observations, to which our locomotion controller can respond appropriately, our task-level terrain estimation module focuses on semantic features such as color and texture. These features are highlighted by the RGB channel, specifically pertaining to terrains such as grass, snow, and cement. To correspond the RGB observation into real-world coordinates, we apply the perspective transformation on $\bm{I}_{\textrm{rgb}}$ to map the pixels to $\bm{M}_{\textrm{BEV}}$, which has corresponding coordinates as $\bm{M}_{T}$. In subsequent processing, we discretize $\bm{M}_{\textrm{BEV}}$ into patches (Fig. 4) and assign the same difficulty within each patch.
To accomplish vision-based terrain estimation, the robot first identifies the observed terrain from $\bm{M}_{\textrm{BEV}}$. We achieve this by training a terrain classification network ${\pi_{\textrm{terrain}}}$ from offline collected data $\bm{D}_{\textrm{terrain}}$. Considering that such inference falls short when faced with unfamiliar terrains, we calculate the predictive entropy to approximate the uncertainty $U$ of the inference at current observation [36]. Since the approximation requires Monte Carlo simulation, we perform $K$ stochastic forward inferences with different data augmentations at each timestep, with $P_{i}^{k}$ denoting the predicted probability for label $i$ at trail $k$. We estimate the expectation as the average of the predicted probability $P_{i}=\sum_{k}P_{i}^{k}/K$. Then the predicted explicit terrain names $T$ and corresponding uncertainty $U$ for the current patch can be calculated as:
For safe navigation, terrain traversability signifies the robot’s preference or stability in traversing a particular terrain, which has been explored from either self-supervision [11] or human demonstrations [24]. We leverage experience from these efforts to establish the terrain traversability (Fig. 4) as the operator preference. As we demonstrated in Fig. 3, such prior knowledge-based terrain traversability aligns with the proprioception evaluations.
Figure 4: The proposed terrain estimator incorporates a visual estimator and online corrections. The visual estimator infers the terrain name $T$ from the BEV map with its uncertainty ${U}$, and the corresponding terrain traversability $\bm{M}_{T_{O}}$ for $T$ is interpreted based on prior knowledge. The online corrections involve recording the motion evaluations $\bm{M}_{T_{P}}$ by the proprioception advisor and retrieving them from the experience buffer when encountering similar observations. In the figure, $(x_{b},y_{b})$, $(x_{o},y_{o})$, and $(x_{r},y_{r})$ represent the locations of the robot base, observation point, and historical position, respectively.
IV-E Online Terrain traversability Corrections
Through the seamless integration of vision and proprioception, we provide online corrections for the vision-based traversability estimation $\bm{M}_{T_{O}}$ without additional training (Fig. 4). This enables the robot with terrain awareness beyond the limitations of the collected data $\bm{D}_{\textrm{terrain}}$.
here $k_{T_{3}}$ and $S_{0}$ are hyperparameters.
here $k_{T_{4}}$ and $C_{0}$ are hyperparameters.
By integrating the adapted $\bm{M}_{T}$ as the terrain traversability into the path planner, our system demonstrates rapid adaptability to different terrain and visual conditions.
V Evaluations
V-A Experimental Setup
We assess TOP-Nav in both simulation and the real world. Considering the sim2real gap primarily lies in the RGB semantic information, we assume the robot has access to ground-truth terrain traversability in simulation. Therefore, our simulation experiments were designed to emphasize the following evaluations:
  • The improvements of Terrain awareness in navigating challenge environments.
  • The improvements of the proposed Proprioception advisor in navigating the challenge environments.
  • The improvements in locomotion by selecting waypoints with higher traversability.
Relatively, in addition to the evaluations mentioned above, our real-world experiments specifically focus on assessing the enhancements provided by the proposed terrain estimator.
Simulation Settings: Our simulation experiments are conducted within Nvidia Isaac Gym. We create a grid of ${8\times 8}$ independent navigation cells. Each cell is ${5m\times 5m}$ in size, featuring a robot assigned to a point goal navigation task. The robot and point goal is randomly generated, with a minimum initial distance of ${5m}$ within the cell. $\bm{M}_{C}$ is configured with dimensions $(1.2m,1.5m)$ centered around the robot‘s location at ${\bm{p}_{\textrm{base}}^{\textrm{map}}=(0.45m,0.75m)}$. This design prioritizes a light-weight, real-time updated path planner and maintains consistency with the height map used in locomotion. The environments include randomly generated obstacles and terrains. Each cell is equipped with at least one wall obstacle, two column obstacles and the remaining space is divided into ${1m\times 1m}$ sections. For terrain traversability assignments, we uniformly partition the obstacle-free space into difficulty levels [0, 0.25, 0.5, 0.75, 1] and generate terrain with various heights of steps and the intensity of irregular terrain corresponding to the difficulty levels. This difficulty serves as ground-truth walking preferences in simulation. The simulation experiments are conducted $25$ times in each of the $64$ navigation cells.
Figure 5: We establish a parallel navigation evaluation environment in simulation, featuring ${8\times 8}$ independent navigation cells. Each cell consists of randomly generated challenging terrains with distinct traverse difficulty, which is marked by the irregularity and complexity of the terrain. The proposed terrain awareness navigation framework plans an optimal path to navigate challenging terrains. The robot demonstrates the capability to recover from unexpected obstacles or irregular terrains with the proprioception advisor.
Real-World Settings: For real-world evaluations, we conduct outdoor navigation tasks in $5$ different scenarios involving challenge obstacles and various terrains, with each scenario replicated $5$ times for each method under investigation. To obtain a more comprehensive observation in the real world, we configure the cost map $\bm{M}_{C}$ with dimensions of $(3m,3m)$. The location of robot is at ${\bm{p}_{\textrm{base}}^{\textrm{map}}=(0m,1.5m)}$ in real world $\bm{M}_{C}$. The criteria for success include reaching the goal within the specified time constraints, consistent with the simulation setting.
Metrics: We evaluate TOP-Nav with the following metrics:
  • SR (Success Rate): The percentage of successful experiments. We define a success experiment as approaching the point goal into ${0.5m}$ within 20 seconds
  • TD (Terrain Difficulty): The percentage of average traversed terrain costs, TD provides the ground truth traversed terrain costs within each episode.
  • UT (Unstable Time): The percentage of unstable motion states ($\lvert roll\rvert>0.15\text{ or }\lvert pitch\rvert>0.15$) in each episode.
  • VFT (Velocity Tracking Failure): The average percentage of velocity tracking failures ($\|v_{\textrm{lin}}-v_{\textrm{act}}\|>0.2$) in each episode.
  • AEC (Average Energy Consumption): The average energy consumption ($\tau\dot{q}$) [13] in each successful episode.
Among them, UT, VFT and AEC demonstrate walking states while completing navigation tasks in challenging terrains., TD provides a reference terrain traversable cost, measuring the improvements in terrain-aware path planning for gait stability. We calculate the variance across different navigation scenarios to assess the robustness of the proposed method.
V-B Simulation Evaluations
TABLE I: Simulation Results with comparison experiments and ablation study
VP-Nav ${65.62\pm 2.38}$ ${47.18\pm 0.79}$ ${36.18\pm 0.54}$ ${26.09\pm 0.42}$ ${101.80\pm 49.61}$
wo/Terrain ${68.00\pm 2.57}$ ${45.83\pm 0.70}$ ${34.92\pm 0.59}$ ${25.06\pm 0.46}$ ${101.70\pm 46.55}$
wo/Proprioception ${62.75\pm 1.64}$ ${\bm{20.47\pm 0.17}}$ ${32.87\pm 0.48}$ ${26.66\pm 0.57}$ ${\bm{88.17\pm 15.31}}$
Obstacle-Only ${52.81\pm 2.30}$ ${47.79\pm 0.82}$ ${41.45\pm 0.79}$ ${39.13\pm 0.99}$ ${96.82\pm 48.03}$
TOP-Nav ${\bm{73.62\pm 1.21}}$ ${22.65\pm 0.19}$ ${\bm{27.21\pm 0.24}}$ ${\bm{18.14\pm 0.16}}$ ${92.08\pm 24.39}$
Improvements with Terrain Awareness: We illustrate a comprehensive terrain and obstacle map within one nav-cell in Fig. 5, exemplifying that the robot equipped with terrain awareness can plan a path with higher traversability. As demonstrated in Table I, when integrated with terrain awareness, TOP-Nav surpasses the VP-Nav baseline by approximately $8\%$ in success rate (SR). We observe that TOP-Nav and the wo/Proprioception methods achieve a TD of nearly $20\%$, which is half of the TD achieved in methods without terrain awareness. These results indicate that the proposed path planner empowers the robot to select terrains with higher traversability, leading to a significant improvement in the success rate of navigation.
Figure 6: In scene 1-3, the green line represents the trajectory of TOP-Nav, while the red line represents Obstalce-Only. Our demonstrations for scene 4,5 are divided into phases that we provide details of the cost map integration in the figure.
Improvements with Proprioception Advisor: The simulation environments are cluttered with various obstacles and difficult terrains, leading to potential locomotion failures such as getting stuck on irregular terrain surfaces or colliding with unseen obstacles during directional changes (Fig. 5). We demonstrate that the proposed proprioception advisor could address those challenges: In contrast to methods without proprioception, our approach exhibits an approximate $11\%$ enhancement in SR. This is evident in the comparison between TOP-Nav and wo/Proprioception, as well as the comparison between wo/Terrain and Obstacle-Only. Such ablations also indicate that the terrain traversability integration and the proprioception advisor each contribute to performance improvement. Meanwhile, even without terrain awareness, the proposed system outperforms VP-Nav by $2.5\%$, signifying an improvement in our integration of the proprioception advisor compared to existing methods.
Advancements in Locomotion: The metrics of UT and VFT evaluate the locomotion states. We observe that TOP-Nav achieves the lowest UT and VFT, indicating that selecting simpler terrains contributes to locomotion stability. In the absence of the proprioception advisor, wo/Proprioception and Obstacle-Only exhibit a significant increase in VFT due to the robot getting stuck by unexpected obstacles. We evaluate energy consumption with the assumption that when traversing simpler terrain, the robot should exhibit more natural gaits. As a result, TOP-Nav and wo/Proprioception exhibit a $10\%$ reduction in energy consumption compared to methods without terrain awareness.
V-C Real World Evaluations
As shown in Fig. 6, TOP-Nav is evaluated in different environments featured with diverse terrains and obstacles. Among them, the terrains encountered in scenes 1-3 were provided in the terrain classifier, we assess the improvements introduced by the vision-based terrain estimator in these experiments. scenes 4-5 are designed to assess the effectiveness of the online correction module when encountering novel terrains. In scene 4, the terrain classifier does not include high-cost gravel. In scene 5, the robot encounters terrains with no prior knowledge, including a slippery detergent surface. The evaluations in scene 6 demonstrate that with the proposed proprioception advisor, the robot is able to avoid invisible obstacles.
Due to significant velocity estimation errors in the real robot, we exclude VFT in the real-world experiments. The Time metric measures the proportion of time the robot takes to complete the navigation task. ST in scenario 5 assesses the time taken by the robot from encountering a transparent object to completely avoiding the obstacle.
Improvements with the Visual Estimator: The evaluation results for scenes 1-3 are presented in Table II. TOP-Nav successfully completes all 15 trials with superior walking stability and minimal energy consumption. These results affirm the successful deployment of the proposed system on real hardware, allowing the robot to select waypoints on terrains with better traversability and thus execute more natural gaits. In contrast, GA-Nav, trained on RUGD [37], demonstrates notable limitations when confronted with open-world navigation. We observe that the improvements of TOP-Nav over GA-Nav mainly stem from: 1) In scene 1,2, GA-Nav fails to identify the slabbed road as a low-cost terrain from the bush and gravel. 2) In scene 3, GA-Nav incorrectly identifies the lower part of the obstacle as a cement floor, which could be addressed by our integration of obstacles and terrain estimation. Nevertheless, GA-Nav surpasses VP-Nav and Obstacle-Only, which navigates without terrain awareness. With Obstacle-Only, the robot gets stuck by the stones or Grassroots, leading to a higher time consumption. While VP-Nav effectively alerts the robot to locomotion failures that could be brought by such terrains, the absence of terrain awareness prevents the robot from successfully navigating out of these challenging terrains, resulting in inferior locomotion performance.
TABLE II: Evaluation results on scene 1,2,3.
@@ V-C Real World Evaluations
TABLE III: Evaluation results on scene 4 (Novel Terrains).
GA-Nav 5/555{5/5}5 / 5 ${\bm{7.98}}$ 62.5762.57{62.57}62.57 67.4267.42{67.42}67.42
VP-Nav ${4/5}$ 26.4326.43{26.43}26.43 70.3070.30{70.30}70.30 77.9077.90{77.90}77.90
Obstacle-Only 5/555{5/5}5 / 5 42.3342.33{42.33}42.33 72.6072.60{72.60}72.60 70.6470.64{70.64}70.64
TOP-Nav 5/555{5/5}5 / 5 14.3014.30{14.30}14.30 ${\bm{57.46}}$ ${\bm{62.72}}$
Improvements with the Online Correction: To evaluate the efficacy of the online corrections for providing the robot with traversability on novel terrains, we conduct experiments in scene 4 using a terrain classifier trained without gravel. The degraded classifier exhibits an accuracy of 96.30% on the paved road with an average confidence of 0.87, whereas the inference on gravel has an average confidence of 0.43. We provide a navigation example in Fig. 6 and divide it into four phases. In phase 1, despite ${\bm{M}_{P}}$ indicating that the robot is traversing a challenging terrain, the goal map ${\bm{M}_{G}}$ guides the robot to continue moving forward until this information is incorporated into ${\bm{M}_{T_{P}}}$. In phase 2, an online correction is applied as the robot encounters similar gravel terrain, leading to an increased cost brought by ${\bm{M}_{T_{P}}}$ in the current direction. Consequently, the robot turns right to tap on the paved road. In phase 3, the robot avoids an observed obstacle. In phase 4, the historical proprioception continues to indicate the presence of high-cost gravel terrain on the left. The quantitative results are provided in Table III. Despite GA-Nav possessing complete prior knowledge and correctly identifying the paved road in scene 4, our online corrections only lead to a $7\%$ increase in UT, mainly due to the forward movement in phase 1. VP-Nav allows the robot to exit the gravel terrain with the safety advisor, but it can not retain this information. As a result, the target orientation prompts the robot to return to the gravel, causing increased time consumption and instability.
TABLE IV: Evaluation results on scene 5 (Novel slippy terrains).
Method SR $\uparrow$ UT(%) $\downarrow$ AEC $\downarrow$ ST(s) $\downarrow$
VP-Nav ${3/5}$ 2.642.64{2.64}2.64 102.04102.04{102.04}102.04 $\bm{63.05}$
Obstacle-Only ${2/5}$ 3.883.88{3.88}3.88 119.77119.77{119.77}119.77 66.1466.14{66.14}66.14
Sterling ${2/5}$ 3.853.85{3.85}3.85 113.75113.75{113.75}113.75 76.5276.52{76.52}76.52
TOP-Nav 5/555{5/5}5 / 5 ${\bm{1.19}}$ ${\bm{83.40}}$ 67.6267.62{67.62}67.62
Evaluations on Challenge Novel Terrains: In Scene 5, we present a more challenging scenario where the robot encounters terrains without any prior information. A black tarpaulin covered with slippery detergent is laid out in front of the robot, and the target direction is set in this facing direction. To mitigate the disturbances brought by reflections from detergents, the obstacle cost ${\bm{M}_{O}}$ is not considered in this experiment. As demonstrated in Fig. 6, the experiment could be divided into three phases. In phase 1, the robot moving forward along the goal direction, but the slippery terrain causes a decline in motion evaluations, resulting in a higher ${\bm{M}_{P}}$. Simultaneously, the traversability cost of the slippery tarpaulin is recorded along with the visual features. In phase 2, the robot begins to turn aside to avoid the high-cost central area. In phase 3, the cost of ${\bm{M}_{G}}$ guides the robot to return to the original direction, where it can navigate on the wooden floor with better traversability to complete the task. The results presented in Table IV validate the effectiveness of our method in maintaining stability and improving the success rate by avoiding challenging terrain. Among the comparisons, Obstacle-Only keeps moving forward on the slippy surface and exhibit the worst motion performance. We include Sterling in this comparison to demonstrate the importance of online corrections when incorporating proprioception into terrain estimation. Without online corrections, Sterling behaves similarly to Obstacle-Only, unable to adapt to novel terrains and continuing forward. Compared with the result in scene 4, we demonstrate that TOP-Nav can handle novel terrains effectively without relying on operator preferences. This showcases the ability of the system to perform open-world navigation.
TABLE V: Evaluation results on scene 6 (Invisible Obstacles).
Method SR $\uparrow$ UT(%) $\downarrow$ AEC $\downarrow$ ST(s) $\downarrow$
VP-Nav 5/555{5/5}5 / 5 1.451.45{1.45}1.45 42.9742.97{42.97}42.97 3.303.30{3.30}3.30
[…]
TOP-Nav 5/555{5/5}5 / 5 ${\bm{0.79}}$ ${\bm{34.63}}$ ${\bm{1.98}}$
Figure 7: In scene 6, the robot avoids unseen obstacles with the proprioception advisor.
Invisible Obstacles: The proposed proprioception advisor provides close-loop feedback to alert the robot to invisible obstacles. We assess this capability in scene 6, where a glass wall is located along the planned path. In phase 1, the obstacle estimator is oblivious to the presence of the glass wall, and the robot continues moving forward. In phase 2, the robot collides with the glass wall, causing a rapid increase in ${\bm{M}_{P}}$ along the direction of movement. As a result, the robot adjusts its waypoint to steer clear of the high-cost area, successfully avoiding the glass wall. The quantitative results are provided in Table V.
Conclusion
We present TOP-Nav, a legged navigation system that achieves closed-loop integration of visual and proprioception at both task and motion planning levels. Specifically, not only have we incorporated vision-based obstacle and terrain estimation into the navigation framework, but we have also utilized the proprioception advisor to make online corrections on these vision-only estimations. To validate the proposed system, we conducted extensive experiments in simulated environments featuring diverse terrains and obstacles. Furthermore, the navigation system was successfully deployed on a real robot, and quantitative experiments were conducted. These evaluation results underscore the success of our system in achieving open-world navigation, surpassing limitations posed by terrain, obstacles, or prior knowledge. Compared to existing approaches that couple vision and proprioception or navigate with terrain awareness, our system excels not only in achieving a higher success rate in challenge navigation, but also in demonstrating more stable gaits and lower energy consumption. In the future, we aim to extend the application of the proposed system to longer-distance navigation tasks through precise localization from radar. Furthermore, we plan to enhance the task planner with data-sufficient models such as Transformer.

### v3

@@ 7.2 Path Planner Details
For reference, we provide the default values of the hyperparameters used in the path planner in Table 7. Parameters $d_{0},t_{0}$ adjust the weights of $\bm{M}_{T}$ and $\bm{M}_{G}$ as the robot approaches the target and as time progresses. $k_{G_{0}}$ keeps a minimum weight for the robot approaching the target. The listed values remain consistent in simulation; however, considering the various conditions in the real world, we make slight changes to these values in different real-world experiments for better validation of the contribution of this work. For instance, in outdoor navigation experiments, we assign a higher value to $k_{T_{1}}$ to expand the scale of the terrain estimator, which is effective for evaluating the accuracy of the proposed terrain estimator.
The default commanded velcoity $v_{0}$ is set to $0.5m/s$ in both simulation and the real world. The maximum distance $d_{\textrm{max}}$ we considered for computing the obstacle map $\bm{M}_{O}$ is $0.3m$.
Table 7: Hyperparameters in the path planner
Parameter $d_{0}$ $t_{0}$ $k_{T_{1}}$ $k_{T_{2}}$ $k_{G_{0}}$ $k_{G_{1}}$ $k_{G_{2}}$
Default Value $0.5$ $0.5$ $1.0$ $2.0$ $0.1$ $0.4$ $-10$
7.3 Motion Planner Details
The motion planner is implemented following [3, 42]. Both the actor and critic networks within the locomotion policy ${\pi}$ have hidden layer sizes of [512, 256, 128]. The proprioception observation $\bm{o}_{t}^{p}$ includes angular velocity (3), Orientation (2), velocity commands (3), joint positions (12), joint velocities (12), and the last action (12). We store the last 10 steps of $\bm{o}_{t}^{p}$ into the history observation $\bm{o}_{t}^{h}$. The depth encoder learns the exteroception latent features from onboard observation $\bm{I}_{d}$ using a conv-GRU structure. ${\pi}$ is updated using PPO [43] for 20K iterations, the batch size is 160000 divided into 4 mini-batches. The depth encoder is optimized for 5k iterations. The rewards obtained at each step are the sum of the reward functions listed in Table 8.
Table 8: Reward Functions
Target Velocity Tracking1 $min(\overline{v_{act}},v_{\textrm{lin}})$
Target Direction Tracking $e^{-\mid\Delta_{\textrm{yaw}}\mid}$
Orientation Penalty $-(roll^{2}+pitch^{2})$
Hip Joint Position Penalty ${-\|q_{hip}\|}^{2}$
z-direction velocity Penalty ${-\mid v_{z}\mid}^{2}$
Collision Penalty2 $-\sum_{i\in(calf,thigh)}({F_{c}^{i}\geq 0.1N})$
Torques Variation Penalty ${-\|\tau_{t}-\tau_{t-1}\|}$
  • 1
    $\overline{v_{act}}$ is the projection component of $v_{act}$ in the target direction.
  • 2
    $F_{c}^{i}$ is the contact force of calf and thigh indices.
7.4 Proprioception Advisor Details
The default $\pi_{\textrm{critic}}$ requires privileged observations, we train another value function estimation network using observable motion states for the proprioception advisor to be deployed on hardware. The estimated motion evaluation $c_{t}$ is normalized using a sigmoid function. The proprioception advisor is trained with the locomotion policy ${\pi}$, with an exact same network architechture as $\pi_{\textrm{critic}}$. The input only includes proprioception observation $\bm{o}_{t}^{p}$ and historical observation $\bm{o}_{t}^{h}$, updated through the MSE loss with $c_{t}^{\textrm{targ}}$. The estimated value $c_{t}$ is normalized through sigmoid function:
$Norm(c_{t})=\frac{1}{1+e^{2*(2.2-c_{t})}}$ (14)
In practice, and we set a minimum threshold $c_{th}$ for the normalized critic value be considered in the path planner, in simulation $c_{th}=0.8$, while in the real world $c_{th}=0.5$, $Norm(c_{t})$ larger than $c_{th}$ will be set to $1$.
$\bm{M}_{P}$ is calculated with $k_{P}=0.3$, we demonstrate an example of $\bm{M}_{P}$ after the robot collides with a wall (with vision and obstacle detection disabled in this trial) in Fig 10. In this scenario, the normalized $c_{t}$ returns $0.02$, $\bm{M}_{P}$ has a width of $0.75m$ along the moving direction of the robot with costs greater than 0.5.
Figure 10: We demonstrate an example of $\bm{M}_{P}$ when the robot collides with an obstacle.
Generally, the proprioception advisor engages when locomotion states are compromised. To enable recovery from being obstructed by the unexpected terrains, we have designed the following recovery strategy in simulation: when $Norm(c_{t})$ falls below a specific threshold (0.5), we initially assign a backward velocity command of 0.5 m/s for 0.5 seconds. Next, to address potential terrain-related constraints that may prevent the robot from reaching planned waypoints at its default velocity, we reduce the frequency of the path planner to 0.5 Hz until the robot reaches the waypoint or makes a directional change of 0.3 radians.
7.5 Terrain Estimator Details
For prior data collection of the terrain estimator, we teleoperate the robot to walk for 5 minutes on each of the concerned terrains. We capture the motion evaluations $c_{t}$ as well as the first-view observations during these demonstrations. We convert the RGB observations into BEV maps and dividing them into patches. The collected data is labelled based on the corresponding terrain from which it was acquired. The classification network is implemented using the MobileNet backbone and trained on a Nvidia GTX 2080 Ti for 6 hours. The proposed pipeline provides a light-weight inference model of 8 MB to be deployed for onboard computation. During deployment, to approximate the uncertainty $U$ with predictive entropy, We perform $K=8$ stochastic forward inferences with different data augmentations as Monte Carlo simulation, We estimate the expectation as the average of the predicted probability $P_{i}=\sum_{k}P_{i}^{k}/K$.
Based on the collected motion evaluations, we compute the average $\bm{T}_{P}$ observed during walks on each terrain. These walking experiences provide reference values for assigning the terrain costs defined by the operator. Table 9 presents the validation accuracy of the terrain classifier, along with the reference terrain cost and the average collected $\bm{T}_{P}$ for each terrain. Since the terrains encountered in the real world did not exhibit significant differences in motion evaluations, we did not strictly set the reference cost based on linear correlation.
Table 9: Terrain Cost and Classify Accuracy
Name Bush Brick Snow Gravel Slab Cement Paved Grass
@@ 7.5 Terrain Estimator Details
$\bm{T}_{P}$ 0.481 0.523 0.445 0.507 0.581 0.580 0.580 0.572
Cost 0.7 0.6 0.9 0.7 0.0 0.0 0.0 0.2
The hyperparameters used for online corrections are fine-tuned based on the distribution of the learned latent features $\bm{L}$ from the pre-collected data. Figure 11-(b) depicts a t-SNE visualization of the features inferred by the terrain classifier, involving data both within and outside of $\bm{D}_{\textrm{terrain}}$. We observe that $\pi_{\textrm{terrain}}$ can learn distinctive features for specific terrains, showcasing unique clustering in the latent space, even for previously unseen terrain types. The average cosine similarity $S_{(o,r)}$ within each terrain class is $0.95$, we set the middle point $S_{0}$ to be $0.85$ and $k_{T_{3}}$ to be $15$. This allows the terrain cost derived from historical experience to decrease rapidly when the similarity falls below $0.8$.
Figure 11: (a) We illustrate the distribution of samples across different confidence levels, with the training of the terrain classifier, the confidence values for known terrains predominantly clustered beyond 0.9, while samples of novel terrains remained dispersed across the confidence axis. (b)T-SNE visualization of the features inferred by the terrain classifier.
To demonstrate the effectiveness of computing confidence in deciding whether the observed terrains is outside of $\bm{D}_{\textrm{terrain}}$, we illustrate the distribution of samples across different confidence levels in Fig 11-(a). After 200 steps of training, the confidence values for known terrains predominantly clustered beyond 0.9, while samples of novel terrains remained dispersed across the confidence axis. The confidence is used to adjust the weight of online correct ions in terrain estimation. We set $k_{T_{4}}=20$ and $C_{0}=0.9$ in the experiments.
7.6 Hardware Details
We implement TOP-Nav on the Unitree-Go1 and Unitree-Go2 quadruped robot, with a weight of approximately $12kg$ and dimensions of $645mm\times 280mm$. The robot is equipped with 12 brushless motors, each capable of producing a torque of $35.5Nm$. The perception module is equipped with a RealSense D435i camera, offering simultaneous depth and RGB channels. All computations are processed onboard with an NVIDIA Jetson NX asynchronously, i.e. the motion planner operates at a fixed frequency of $50Hz$ and receives the latest depth latent $\bm{\ell}$ inference from the perception module. The path planner operates at a frequency of ${3Hz}$.